### v1

# Generic orbit recovery from invariants of very low degree

Abstract.
Motivated by the multi-reference alignment (MRA) problem and questions in equivariant neural networks we study the problem of recovering the generic orbit in a representation of a finite group from invariant tensors of degree at most three. We explore the similarities and differences between the descriptive power of low degree polynomial and unitary invariant tensors and provide evidence that in many cases of interest they have similar descriptive power. In particular we prove that for the regular representation of a finite group, polynomial invariants of degree at most three separate generic orbits answering a question posed in [4]. This complements a previously known result for unitary invariants [29]. We also investigate these questions for subregular representations of finite groups and prove that for the defining representation of the dihedral group, polynomial invariants of degree at most three separate generic orbits. This answers a question posed in [10] and it implies that the sample complexity of the corresponding MRA problem is $\sim\sigma^{6}$. On the other hand we also show that for the groups $D_{n}$ and $A_{4}$ generic orbits in the complete multiplicity-free representation cannot be separated by invariants of degree at most three.
1. Introduction
Let $V$ be a representation of a reductive group $G$. A classical problem in invariant theory is to give bounds on the number and degrees of generators for the invariant ring $k[V]^{G}$. A more recent problem, introduced by Derksen and Kemper [14], is to find similar bounds for separating invariants - that is collections of invariants that separates all orbits [16, 15]. Motivated by a broad range of applications, we consider the related problem of determining lower bounds on the degrees of invariants necessary to separate generic orbits in a representation of a compact group [4, 19, 11]. This problem has two important applications. The first is the multi-reference alignment (MRA) problem, which arises in cryo-electron miscroscopy. The second is the problem of constructing efficient equivariant neural networks. In both applications it is imperative to separate generic orbits with invariants of the lowest possible degree.
Remarkably, the degrees of invariants needed to separate generic orbits can be significantly lower than the degrees of invariants needed to separate all orbits. For example if $G=\mathbb{Z}_{p}$ with $p$ prime acting by cyclic shift on $\mathbb{C}^{p}$ then invariants of degree $p$ are needed to separate all orbits [15], while it is known that invariants of degree at most three separate generic orbits [4].
Previous papers studied the generic orbit separation problem for band-limited functions in $L^{2}(G)$, where $G$ is a connected compact group [19], and finite abelian groups [4, 11]. This paper is part of a program for understanding when invariants of degree at most three can separate generic orbits. This is the smallest degree $d$ for which we can expect to find large families of representations for which generic orbits are separated by invariants of degree at most $d$ [9]. The main focus of this paper is on representations of finite groups which we view as a special case of compact groups. Because we work with compact, as opposed to reductive groups, we consider both polynomial invariants (which are the same as the polynomial invariants of the reductive complexification of the compact group) as well as unitary invariants such as the power spectrum in degree two and the bispectrum in degree three. (See Section 2 for definitions.) These invariants coincide for real vectors, but for complex vectors they carry different information and it may be possible to separate generic orbits with one sets of invariants but not the other.
1.1. Results
1.1.1. Regular representations of finite groups.
Our first result, Theorem 4.1, states that if $G$ is finite group and $V$ is any representation containing a copy of the regular representation then generic orbits are separated by polynomial invariants of degree at most three. This complements a well-known result in the literature [29] states that for any compact group $G$, the bispectrum (which is equivalent the collection of unitary invariants of degree three) separates generic orbits in the regular representation $L^{2}(G)$. In the case when $G$ is finite this implies that the generic orbit in the regular representation is determined by both the polynomial and unitary invariants of degree at most three. This was previously known for finite abelian groups [4].
1.1.2. Subregular representations and the standard representation of the dihedral group.
We then turn our attention to subregular representations of finite groups. Our main positive result is Theorem 4.1 which states that for the dihedral group $D_{n}$ acting on $\mathbb{C}^{n}$ by cyclic shifts and reflection, the polynomial invariants of degree at most three separate generic orbits. However, unlike for the regular representation, we do not know if the corresponding unitary invariants also separate generic orbits. Our proof of Theorem 4.1 involves several steps. We first consider the separation problem for real orbits (Section  5) and then use algebro-geometric techniques to bootstrap from the real to complex orbit separation (Section 6). Our proof of Theorem 4.1 for real vectors takes advantage of the fact that in the Fourier domain the standard real representation of $D_{n}$ is the restriction of a band-limited representation of the orthogonal group $O(2)$, which we view as the ‘continuous’ dihedral group. We then prove Theorem 5.1, a result of independent interest, stating that for this real representation the $O(2)$-invariants of degree at most three separate generic orbits. The proof Theorem 4.1 for real vectors is completed by proving that if two vectors in an $O(2)$-orbit have the same dihedral invariants then they lie in a common $D_{n}$ orbit.
1.1.3. The complete multiplicity-free representation and other subregular representations of finite groups.
The standard $n$-dimensional representation of the dihedral group $D_{n}$ is a sum with multiplicity one of all but one ($n$ odd) or two ($n$ even) of the irreducible representations of $D_{n}$. We define the complete multiplicity free representation of a finite group to be the sum of all irreducibles each taken with multiplicity one. Interestingly we show in Corollary 4.4 that generic orbits in the complete multiplicity-free representation of $D_{n}$ cannot be separated by invariants (polynomial or unitary) of degree three or less. This phenomenon occurs for other finite groups such as $A_{4}$ and a question for further work is to determine the minimum positive integer $r$ such that invariants of degree at most three separate generic orbits in the representation $V=\oplus_{W\in Irrep(G)}W^{r}$. Our first result implies that this is true if $r\geq\max_{W\in Irrep(G)}\dim W$. In Section 4.3 we analyze this problem for the permutation action of the symmetric group $S_{n}$ on $n\times d$ matrices, since this representation naturally arises in deep learning [3, 28, 30, 17].
1.2. Connection to multi-reference alignment and cryo-EM
Given a representation $V$ of a compact group $G$, the multi-reference alignment (MRA) problem is that of recovering the orbit of a signal vector $x\in V$ from its noisy translates by unknown random group elements $g_{i}$
(1) $y_{i}=g_{i}\cdot x+\epsilon_{i}$
where the noise vectors $\epsilon_{i}$ are taken from a Gaussian distribution $N(0,\sigma^{2}I)$ which is independent of the group elements $g_{i}$. This model was studied in [5] for the regular representation of $\mathbb{Z}_{N}$ and in [6] it was proposed as an abstract version of the cryo-EM signal reconstruction problem. The MRA problem for the regular representation of $\mathbb{Z}_{n}$ has been extensively studied in recent years, with both uniform and non-uniform distributions being considered [24, 8, 1, 7]. Other models include the dihedral group with non-uniform distribution [10] and rotation groups acting with uniform distributions on spaces of band-limited functions in $\mathbb{R}^{2}$ and $\mathbb{R}^{3}$ [6, 23, 20, 19].
When the signal-to-noise ratio is extremely low, as is the case for cryo-EM measurements, there is no way to estimate the unknown group elements, but it can be shown that the moments of the unknown signal can be accurately approximated [2]. When the distribution of random group elements is governed by a uniform distribution the moments are invariant tensors and the MRA problem reduces to the problem of recovering an orbit from its invariants. However, the sample complexity (the minimal number of measurements necessary for accurate approximation) grows exponentially in the degrees of the moments, so to efficiently solve the MRA problem it is necessary to recover almost all signals from invariants of the lowest possible degree.
As a corollary of Theorem 4.1 we obtain the following result about the sample complexity of the MRA problem for the standard representation $\mathbb{C}^{n}$ of the dihedral group $D_{n}$.
Corollary 1.1.
[…]
1.3. Connection to deep learning
In machine learning it is desirable to build neural networks that reflect the intrinsic symmetry of the data such as point clouds. If $G$ is the symmetry group of the data then we want to build the network from $G$-equivariant functions and a basic model of an equivariant neural network [13, 22] is a sequence of maps
$\mathbb{R}^{n_{0}}\stackrel{{\scriptstyle A_{1}}}{{\to}}\mathbb{R}^{n_{1}}\stackrel{{\scriptstyle\sigma_{b_{1}}}}{{\to}}\mathbb{R}^{n_{2}}\ldots\stackrel{{\scriptstyle\sigma_{b_{k-1}}}}{{\to}}\mathbb{R}^{n_{k-1}}\stackrel{{\scriptstyle A_{k}}}{{\to}}\mathbb{R}^{n_{k}}$
where each $\mathbb{R}^{n_{i}}$ is a representation of $G$, the $A_{i}$ are $G$-equivariant linear transformations and the $\sigma_{b_{i}}$ are non-linear maps. A difficulty with this model is that there are in general relatively few $G$-equivariant linear maps between representations, so an equivariant network built this way may not be sufficiently expressive. One approach to circumvent this problem is to use invariant polynomial maps rather than linear maps [21, 18, 12]. However, because the cost of computing invariants grows exponentially in the degree, it is important to identify representations for which invariants of low degree separate almost all orbits.
2. Unitary and polynomial invariant tensors
Consider the orthogonal or unitary action of a compact group $G$ on a vector space $V$ which we identify as $\mathbb{R}^{n}$ or $\mathbb{C}^{n}$ depending on context. In this paper we work with two slightly different invariant tensors that arise from the MRA problem and other questions in signal processing.
Definition 2.1 (Moment tensor).
If $V$ is a unitary representation of a compact group $G$ and $x\in V$ then we define the $d$-th moment tensor as
(2) $M^{G}_{d}(x)=\int_{G}(g\cdot x)^{\bigotimes d-1}\otimes(\overline{g\cdot x)}\;dg$
where $dg$ is the uniform Haar measure on $G$.
Definition 2.2 (Invariant tensor).
If $V$ a representation of a compact group $G$ and $x\in V$ we define the $d$-th invariant tensor as
[…]
Note that these tensors agree on real vectors and as such they only differ for complex signals.
When $G$ is finite the moment and invariant tensors are just the summations
(4) $M^{G}_{d}(x)=\sum_{g\in G}(g\cdot x)^{\bigotimes d-1}\otimes(\overline{g\cdot x})$
and
(5) $T^{G}_{d}(x)=\sum_{g\in G}(g\cdot x)^{\bigotimes d}$
By definition, both the moment and invariant tensors are invariants; i.e. for any $h\in G$, $T_{d}^{G}(h\cdot x)=T_{d}^{G}(x)$ and $M_{d}^{G}(h\cdot x)=M_{d}^{G}(x)$. In particular we may view the $d$-th moment tensor as a map $V\to(V^{\otimes d-1}\otimes V^{*})^{G}=\operatorname{Hom}_{G}(V^{\otimes d-1},V)$ and the $d$-th invariant tensor as a degree-$d$ polynomial map $V\to(V^{\otimes d})^{G}$.
The coefficients of $T_{d}^{G}(x)$ are exactly the $G$-invariant polynomials of degree $d$ in $x$. The coefficients of $M_{d}^{G}(x)$ are also $G$-invariant functions but are not polynomials due to the presence of the conjugated terms. We will refer to these non-polynomial invariants occurring in the moment tensor as unitary invariants. We will illustrate the difference with the following example.
Example 2.3.
Let $\mathbb{C}^{n}$ be the standard representation of the cyclic group $\mathbb{Z}_{n}$ represented in the Fourier basis. In other words an element $\ell\in\mathbb{Z}_{n}$ acts by the rule
$\ell\cdot(x_{0},\ldots,x_{n-1})=(x_{0},e^{2\pi\iota\ell}x_{1},\ldots e^{2\pi\iota(n-1)\ell}x_{n-1})$
Because the action is diagonalized all invariants are monomials. There is a single unitary and polynomial invariant of degree one, namely the function $x_{0}$. The degree two polynomial invariants are $\{x_{i}x_{n-i}\}$ whereas the degree two unitary invariants are $\{x_{i}\overline{x_{i}}\}$. In signal processing the set of degree two unitary invariants is called the power spectrum. The degree three polynomial invariants are the monomials $x_{i}x_{j}x_{n-i-j}$ where all indices are taken modulo $n$. Likewise the degree-three unitary invariants are the monomials $x_{i}x_{j}\overline{x_{i+j}}$. In signal processing the collection of degree-three unitary invariants is called the bispectrum.
Note that if $x=(x_{0},\ldots,x_{n})$ is the Fourier transform of a real vector then $x_{i}=\overline{x_{n-i}}$ so the polynomial and unitary invariants agree on these vectors.
Since $\mathbb{C}^{n}$ with our chosen action of $\mathbb{Z}_{n}$ is the regular representation, we know, by Tannaka-Krein duality [29] that the unitary invariants of degree at most three separate generic orbits. Likewise, our Theorem 3.1 implies that the polynomial invariants of degree at most three also separate generic orbits.
Example 2.4.
Now let let $S^{1}=U(1)$ act on $\mathbb{C}^{2}=\mathbb{C}_{1}+\mathbb{C}_{2}$ by the weights 1 and 2 i.e. $\theta\cdot x_{1}=e^{i\theta}x_{1}$ and $\theta\cdot x_{2}=e^{2i\theta}x_{2}$. There are no non-constant polynomial invariants and hence recovery from these invariants is not possible. However, the unitary invariants $x_{1}\overline{x_{1}},x_{2}\overline{x_{2}},x_{1}^{2}\overline{x_{2}}$ separate any $S^{1}$ orbits satisfying $x_{1}x_{2}\neq 0$.
Remark 2.5.
Note that the representation in Example 2.3 is defined over the reals, since it is the regular representation of a finite group. By contrast, the representation we consider in Example 2.4 is not defined over the reals. A natural, but open question, which arises from this work is whether for real representations of finite groups the separating power of algebraic and unitary invariants on complex vectors is the same. This is discussed further in the next section.
The following proposition illustrates another difference between the polynomial and unitary invariants.
Proposition 2.6.
If $V$ contains a copy of the trivial representation then for any $x$ whose projection to the trivial component of $V$ is non-zero then the unitary invariants $M^{G}_{1}(x)$, $M^{G}_{2}(x)$ are determined by $M^{G}_{3}(x)$.
Proof.
Let $f$ be a unitary invariant of degree 2 occurring in the moment tensor $M^{G}_{2}(x)$ and let $x_{0}\in V_{0}$ be the non-zero projection of $x\in V$ to a copy of the trivial representation $V_{0}\subset V$. The unitary invariant $x_{0}f$ is a coefficient in the tensor $M^{G}_{3}(x)$ and hence if the value of $x_{0}$ can be determined then the value of $f$ is determined. Note that $x_{0}x_{0}\overline{x_{0}}=x_{0}|x_{0}|$ is a also a coefficient in the tensor $M^{G}_{3}(x)$ and uniquely determines the value of $x_{0}$. ∎
Remark 2.7.
We cannot repeat this argument for polynomial invariants due to the fact that $x_{0}^{3}$ only determines the value of $x_{0}$ up to a cube root of unity.
2.1. Polynomial invariants and orbit separation
A basic result in the theory of compact Lie groups states that any compact group $G$ is the maximal compact subgroup of a corresponding complex reductive algebraic group $G_{\mathbb{C}}$. Moreover, if $V$ is a unitary representation of $G$ then the polynomial invariants of the compact Lie group $G$ are the same as the polynomial invariants of the algebraic group $G_{\mathbb{C}}$. For example if $G=\operatorname{SU}(n)$ then the corresponding complex algebraic group is $G_{\mathbb{C}}=\operatorname{SL}(n)$. If $V$ is a unitary representation of $\operatorname{SU}(n)$ and $x=(x_{1},\ldots,x_{n})\in V$ then any $\operatorname{SU}(n)$-invariant polynomial $f(x_{1},\ldots,x_{n})$ is also $\operatorname{SL}_{n}$-invariant.
Because of this limitation, complex representations of positive-dimensional compact groups will never have generic separation of orbits from invariant polynomials. However, the unitary invariants can separate orbits in many cases [19]. Here we focus on finite groups, which are both compact and reductive, and thus we will not encounter this difficulty.
As the next example shows, a set of real invariants which separate generic real orbits need not separate complex orbits. The main goal of this section is to give a criterion (Proposition 2.15) for determining when such a collection of real invariants also separate generic complex orbits.
Example 2.8.
Consider the action of $\mathbb{Z}_{3}$ on $\mathbb{C}^{3}$ by cyclic shifts. Under this action $\mathbb{C}^{3}$ decomposes into two complementary $\mathbb{Z}_{3}$-invariant subspaces the diagonal $\Delta=\{(a,a,a)|a\in\mathbb{C}\}$ and its orthogonal complement $V=\{(x,y,z)|x+y+z=0\}$. With respect to this decomposition the invariant polynomial $x+y+z$ determines the component of a vector in $\Delta$ and the degree-two and degree-three invariants
(6) $\left\{\begin{array}[]{l}x^{2}+y^{2}+z^{2}-(xy+xz+yz),\\ x^{3}+y^{3}+z^{3}+3(xy^{2}+x^{2}z+yz^{2})-6(x^{2}y+y^{2}z+xz^{2})+6xyz,\\ x^{3}+y^{3}+z^{3}-6(xy^{2}+x^{2}z+yz^{2})+3(x^{2}y+y^{2}z+xz^{2})+6xyz\end{array}\right\}$
separate all orbits in $V$ and vanish on $\Delta$.
Since a real number is uniquely determined by its cube we can replace the degree-one invariant $x+y+z$ with $(x+y+z)^{3}$ and the set of invariants in degrees two and three
(7) $\left\{\begin{array}[]{l}(x+y+z)^{3},x^{2}+y^{2}+z^{2}-(xy+xz+yz),\\ x^{3}+y^{3}+z^{3}+3(xy^{2}+x^{2}z+yz^{2})-6(x^{2}y+y^{2}z+xz^{2})+6xyz,\\ x^{3}+y^{3}+z^{3}-6(xy^{2}+x^{2}z+yz^{2})+3(x^{2}y+y^{2}z+xz^{2})+6xyz\end{array}\right\}$
separates all real orbits but cannot separate any complex orbit for which the the $\Delta$ component is non-zero.
Although the invariants above do not generate the field of rational functions, they still form a transcendence basis. This implies that although the invariants may not separate generic complex orbits, they must determine the complex orbits up to a finite list of possibilities.
As observed in [4] a fundamental theorem of Rosenlicht [25, Theorem 2] implies the following result.
Theorem 2.9 (Rosenlicht).
Let $V$ be a real or complex representation a finite group $G$ and let $U\subset k[V]^{G}$ be a finite set of polynomial invariants where $k=\mathbb{R}$ or $k=\mathbb{C}$. If $U$ separates generic complex orbits, then $\operatorname{Frac}(\mathbb{C}[U])=\operatorname{Frac}(\mathbb{C}[V]^{G})$.
Remark 2.10.
The result of Rosenlicht holds for any complex reductive group and identifies $\operatorname{Frac}(\mathbb{C}[U])$ with the field of invariants $\mathbb{C}(V)^{G}$. For a general representation $V$ of a reductive group $G$, $\operatorname{Frac}(\mathbb{C}[V]^{G})$ may be a proper subfield of $\mathbb{C}(V)^{G}$. For example if $G=\mathbb{C}^{\times}$ acts on $\mathbb{C}^{2}$ with weights $(1,1)$ then the invariant ring $\mathbb{C}[V]^{G}$ consists only of constants but the field of $G$-invariant rational functions is the field $\mathbb{C}(x/y)$. However, if the identity component of $G$ has no non-trivial characters, as is the case for finite groups, then $\mathbb{C}(V)^{G}=\operatorname{Frac}(\mathbb{C}[V]^{G})$.
Definition 2.11.
[2] Let $G$ be a compact group and $k=\mathbb{R}$ or $k=\mathbb{C}$ and let $U\subset k[V]^{G}$ be a set of polynomial invariants (respectively unitary invariants). We say $U$ list separates generic orbits of $V$ if there is a $G$-invariant Zariski dense subset $W\subset V$ such that for any point $x\in W$ there are only finitely many orbits $\Theta_{1},\ldots,\Theta_{t}$ such that if $p(y)=p(x)$ for all $p\in U$ then $y\in\Theta_{\ell}$ for some $\ell$.
For finite groups, the following proposition, which follows from [4, Theorem 3.15], characterizes when a set of polynomial invariants list separates generic orbits.
Proposition 2.12.
Let $V$ be a real or complex representation a finite group $G$ and let $U\subset k[V]^{G}$ be a finite set of polynomial invariants where $k=\mathbb{R}$ or $k=\mathbb{C}$. Then $U$ list separates generic orbits if and only if $\operatorname{trdeg}_{k}\operatorname{Frac}(k[U])=\operatorname{trdeg}_{k}\operatorname{Frac}(k[V]^{G})$ and in this case the extension of fields $\operatorname{Frac}(k[U])\subset k(V)$ is finite.
Note that if $U$ is a collection of real invariant polynomials then $\operatorname{trdeg}_{\mathbb{R}}\operatorname{Frac}(R[U])=\operatorname{trdeg}_{\mathbb{C}}\operatorname{Frac}(\mathbb{C}[U])$. It follows that if $U\subset\mathbb{R}[V]^{G}$ list separates generic real orbits then it also list separates generic complex orbits, although list sizes may be different. As a result we obtain the following useful corollary comparing the separation properties of the moment and invariant tensors.
Corollary 2.13.
Suppose $V$ is a representation of a finite group $G$. If the moment tensors $\{M_{\ell}^{G}\}_{\ell\leq d}$ list separate generic complex orbits then the corresponding invariant tensors $\{T^{G}_{\ell}\}_{\ell\leq d}$ also list separate generic complex orbits - although the list sizes may be differ.
Proof.
This follows because the moment and invariant tensors agree on real vectors, so the hypothesis that the moment tensors list separate generic complex orbits also implies that the invariant tensors list separate generic real orbits and hence generic complex orbits. ∎
Remark 2.14.
If a collection of invariant tensors list separates generic complex orbits we cannot use the algebraic argument of Corollary 2.13 to conclude that the corresponding moment tensors also list separate generic complex orbits.
We now come to our criterion for determining when a collection of real invariants which separate generic real orbits also separate generic complex orbits.
Proposition 2.15.
Let $G$ be a finite group acting on a real vector space $V=\mathbb{R}^{n}$, and let $U\subset\mathbb{R}[V]^{G}$ be a set of invariants which separates the generic real orbits of $V$. Assume in addition the following property holds:
(*) For generic $x\in\mathbb{R}^{n}$ any $y\in\mathbb{C}^{n}=V\otimes\mathbb{C}$ with the property that $p(y)=p(x)$ for all $p\in U$ implies that $y\in\mathbb{R}^{n}$
Then the invariants $U$ separate generic orbits in $V\otimes\mathbb{C}$ and consequently $\operatorname{Frac}(\mathbb{C}[U])=\operatorname{Frac}(\mathbb{C}[V]^{G})$
Proposition 2.15 follows from the following lemma.
Lemma 2.16.
Let $V=\mathbb{R}^{n}$ be a real representation of a finite group $G$ and let $U\subset\mathbb{R}[V]^{G}$ be a collection of invariants such that $\operatorname{trdeg}_{\mathbb{C}}\operatorname{Frac}(\mathbb{C}[V]^{G})=\operatorname{trdeg}_{\mathbb{C}}\operatorname{Frac}(\mathbb{C}[U])$ and the map $\mathbb{C}^{n}\xrightarrow{\phi}\operatorname{Spec}\mathbb{C}[U]$ has the property that for a generic point $x\in\mathbb{R}^{n}\subset\mathbb{C}^{n}$ the saturation $\phi^{-1}\phi(x)$ consists of $r=[\operatorname{Frac}(\mathbb{C}[V]):\operatorname{Frac}(\mathbb{C}[V]^{G})]$ real points. Then $\phi$ is generically finite of degree $r$ and $\operatorname{Frac}\mathbb{C}[U]=\operatorname{Frac}(\mathbb{C}[V]^{G})$.
Proof.
We note that because $U$ is a transcendence basis, the map $\mathbb{C}^{n}\xrightarrow{\phi}V$ is generically finite of generic degree $m$. In particular there is a dense open set $W\subset\mathbb{C}^{n}$ such that for all $z\in W$ the saturation $\phi^{-1}\phi(z)$ consists of $m$ points. On the other hand we know by hypothesis that for a generic point in $y\in\mathbb{R}^{n}$ $\phi^{-1}\phi(y)$ consists of $r$ real points. Since $\mathbb{R}^{n}$ is Zariski dense in $\mathbb{C}^{n}$ it follows that there is a Zariski dense set in $\mathbb{C}^{n}$ on which the fibers have size $r$. Hence the degree of the generically finite morphism $\mathbb{C}^{n}\to\operatorname{Spec}\mathbb{C}[U]$ must be $r=[\operatorname{Frac}(\mathbb{C}[V]):\operatorname{Frac}(\mathbb{C}[V]^{G})]$. ∎
3. The regular representation of a finite group
It was established in [29] that if $G$ is a compact group and $V=L^{2}(G)$ is the regular representation and $f\in L^{2}(G)$ is a function whose Fourier coefficients are all invertible then the orbit of $f$ is determined from the third moment tensor $M_{3}^{G}(f)$. The proof uses Tannaka-Krein Duality along with other results in abstract harmonic analysis. When $G$ is finite the question as to whether a similar uniqueness result holds for the invariant tensor $T_{3}^{G}$ was posed in [4]. In [4] the authors prove this when $G$ is a finite abelian group. For $G$ an arbitrary finite group they use Jennrich’s algorithm for tensor decomposition to give an algorithmic proof that the orbit of any real valued function $f\in L^{2}(G)$ can be recovered from the invariants of degree at most three - provided that the orbit of $f$ consists of linearly independent functions. (Note that for a real valued function the unitary and polynomial invariants coincide.) The main theorem of this section is an algorithmic proof that if $V$ is a real or complex representation of a finite group $G$ then the orbit of any vector $x\in V$ can be recovered from the first three invariant tensors provided that the orbit $G\cdot x$ consists of linearly independent vectors. By an observation of Satriano [4], a representation contains orbits consisting of linearly independent vectors if and only if it contains a copy of the regular representation.
Theorem 3.1.
Let $V$ a representation of a finite group over $\mathbb{C}$. If the $G$-orbit of a vector $x$ consists of linearly independent vectors, then it is uniquely determined from the invariant tensor $T_{3}(x)$ and one of the invariants $T_{1}(x)$ or $T_{2}(x)$.
Proof.
Given $x\in G$ such that its orbit consists of linearly independent vectors let $Tx=\sum_{g\in G}gx\otimes gx\otimes gx$ be its third invarian tensor. By assumption we know that the vectors $\{gx\}_{g\in G}$ are linearly independent. Let $W$ be the span of these vectors. Since $W$ has a $G$-invariant basis, $W$ is a $G$-invariant subspace of $V$ isomorphic to the regular representation $R(G)$. The invariant tensors $T^{\ell}x$ necessarily lies in $\operatorname{Sym}^{\ell}W$, so it we can replace $V$ with $W$; and hence we can assume that $V$ is the regular representation.
Enumerate the elements of $G$ as $g_{0},\ldots,g_{n-1}$ where $g_{0}=e$ and let $u_{i}=g_{i}x$. Fix two linearly independent elements $a,b\in V^{*}$ and consider the contractions $T_{a}=\sum_{i=0}^{n-1}\langle a,u_{i}\rangle u_{i}\otimes u_{i}$ and $T_{b}=\sum_{i=0}^{n-1}\langle b,u_{i}\rangle u_{i}\otimes u_{i}$ in $Sym^{2}W$. Choosing an ordered basis for $V$ we can view $T_{a}$, $T_{b}$ as symmetric $n\times n$-matrices.
Let $D_{a}=\mathrm{diag}(\langle a,u_{i}\rangle)$ and $D_{b}=\mathrm{diag}(\langle b,u_{i}\rangle)$ then $T_{a}=UD_{a}U^{T}$ and $T_{b}=UD_{b}U^{T}$ where $U=(u_{1},\ldots,u_{n})$ is the matrix whose columns are $\{u_{i}\}$. For a generic choice of $a,b\in V^{*}$ the inner products $\langle a,u_{i}\rangle$ and $\langle b,u_{i}\rangle$ are all non-zero, and the products $\langle a,u_{i}\rangle\langle b,u_{i}\rangle^{-1}$ are distinct. Thus the matrix $T_{a}T_{b}^{-1}=UD_{a}D_{b}^{-1}U^{-1}$ which can be determined from the third invariant tensor has distinct eigenvalues, and hence one-dimensional eigenspaces. To determine the $u_{i}$ we can use the following strategy. Let $u$ be any eigenvector of $T_{a}T_{b}^{-1}$. Then $u=chx$ for some $h\in G$ and non-zero constant $c\in\mathbb{C}$. Then $\{gu\}_{g\in G}=\{cu_{i}\}$. Hence $\sum_{g\in G}gu^{\otimes 3}=c^{3}T_{3}(x)$. Since $T_{3}(x)$ is known, we determine $c^{3}$. If in addition, either $T_{2}(x)$ or $T_{1}(x)$ are known then we can also determine $c^{2}$ or $c$ by comparing $\sum gu^{\otimes 2}$ with $T_{2}(x)$ or $\sum gu$ with $T_{1}(x)$. ∎
Remark 3.2.
The same argument shows that if the orbit of $x$ consists of linearly independent vectors then the third moment $M_{3}(x)$ determines the orbit of $x$. Indeed contracting with generic elements $a,b\in V^{*}$ determines Hermitian matrices $T_{a},T_{b}$ and the eigenspaces of the matrix $T_{a}T_{b}^{-1}$ are all one dimensional and generated by the $u_{i}$. Thus, if $u$ is any eigenvector then $u=chx$ for some constant $c$ and $h\in G$. Comparing $\sum gu^{\otimes 2}\otimes\overline{gu}$ with the known value $M_{3}(x)$ gives $c^{2}\overline{c}=\|c\|c$, from which the unknown constant $c$ can be uniquely determined.
Corollary 3.3.
Let $V$ be a representation of a finite group $G$ which contains a copy of the regular representation and let $U_{\leq 3}$ be the set of polynomial invariants of degree at most three in $\mathbb{C}[V]^{G}$. Then, for a generic $x\in V$, the orbit is determined from the functions in $U_{3}$. In more algebraic language, $\operatorname{Frac}(\mathbb{C}[U_{\leq 3})^{G}=\operatorname{Frac}(\mathbb{C}[V]^{G})$. In particular for $x\in Z(U_{\leq 3})^{c}$ the orbit of $x$ is uniquely determined by the invariants of degree at most three.
Proof of Corollary 3.3.
If $V$ contains a copy of the regular representation, then the generic orbit consists of linearly independent vectors. Thus by Theorem 2.9, we know that
$\operatorname{Frac}(\mathbb{C}[U_{\leq 3}])^{G}=\operatorname{Frac}(\mathbb{C}[V]^{G})$. Hence, every invariant polynomial is expressible as a rational function of the polynomials in $U_{\leq 3}$. In particular, if $x$ is a vector in $Z(U_{\leq 3})^{c}$ then the value of every invariant polynomial in $x$ is determined by the invariants of degree at most three. Since an orbit is uniquely determined by all invariant polynomials, we conclude that $x$ is determined by the invariants of degree at most three. ∎
Example 3.4.
Let $\mathbb{C}^{4}$ be representation of $\mathbb{Z}_{4}$ where $\mathbb{Z}_{4}$ acts by cyclic shifts. The invariants degree 1,2 and 3 in the standard basis are as follows:
(8) $\left\{\begin{array}[]{l}x_{1}+x_{2}+x_{3}+x_{4},x_{1}^{2}+x_{2}^{2}+x_{3}^{2}+x_{4}^{2},x_{1}x_{2}+x_{2}x_{3}+x_{3}x_{4}+x_{1}x_{4},x_{1}x_{3}+x_{2}x_{4},\\ x_{1}^{3}+x_{2}^{3}+x_{3}^{3}+x_{4}^{3},x_{1}^{2}x_{2}+x_{1}x_{4}^{2}+x_{2}^{2}x_{3}+x_{3}^{2}x_{4},x_{1}^{2}x_{3}+x_{1}x_{3}^{2}+x_{2}^{2}x_{4}+x_{2}x_{4}^{2},\\ x_{1}^{2}x_{4}+x_{1}x_{2}^{2}+x_{2}x_{3}^{2}+x_{3}x_{4}^{2},x_{1}x_{2}x_{3}+x_{1}x_{2}x_{4}+x_{1}x_{3}x_{4}+x_{2}x_{3}x_{4}\end{array}\right\}$
We know by Theorem 3.1 or that these 9 invariant polynomials determine the $\mathbb{Z}_{4}$ orbit of a generic $x\in\mathbb{C}^{4}$. However, not all orbits are determined by these invariants. For example if $x=(x_{1},x_{2},x_{3},x_{4})\in\mathbb{C}^{4}$ has vanishing Fourier coefficient $\hat{x}[2]=x_{1}-x_{2}+x_{3}-x_{4}=0$ then the $\mathbb{Z}_{4}$-orbit of $x$ cannot be determined by the invariants (8). This is clear when we write the invariants as monomials in the Fourier basis. In this basis the invariants (8) of degree at most three are $\hat{x}[0],\hat{x}[1]\hat{x}[3],\hat{x}[2]^{2},\hat{x}[1]^{2}\hat{x}[2],\hat{x}[3]^{2}\hat{x}[2]$. If $\hat{x}[2]=0$ then any vector $x^{\prime}$ whose Fourier transform is $(\hat{x}[0],e^{\iota\theta}\hat{x}[1],0,e^{-\iota\theta}\hat{x}[4])$ will have the same invariants of degree at most three. To separate the $\mathbb{Z}_{4}$-orbits of $x$ and $x^{\prime}$ we need to use the degree-four invariant $\hat{x}[1]^{4}$. Note that, since $\mathbb{Z}_{4}$ has order $4$, Noether’s bounds ensures that any two orbits can be separated by invariants of degree at most $4$. In general, for the standard representation of $\mathbb{Z}_{n}$ on $\mathbb{R}^{n}$, we may need invariants up to degree $n$ to separate all orbits, even though generic orbits are separated by invariants of degree at most three by Theorem 3.1.
4. Subregular representations of finite groups
Theorem 3.1 implies that the generic orbit in any super-regular (ie a representation that contains a copy of the regular representation) of a finite group can be recovered from the first three moment or invariant tensors. A natural problem is to understand sub-regular representations of finite groups for which the the generic orbit can be recovered from the first three moments or invariant tensors.
4.1. The standard representation of the dihedral group
Let $D_{n}$ be the dihedral group of order $2n$ with generators $r$ of order $n$ and $s$ of order two. Consider the $n$-dimensional standard representation of $D_{n}$ where the generator $r$ acts by cyclic shifts and the generator $s$ acts by the reflection $s(x_{0},\ldots x_{n-1})=(x_{0},x_{n-1},x_{n-2},\ldots x_{1})$.
Our main positive result for the dihedral group is the following theorem which will be proved in Section 5.
Theorem 4.1.
The first three invariant tensors separate generic complex orbits in the standard representation of $D_{n}$.
Using the same argument as in the proof of Corollary 3.3 we obtain the following corollary.
Corollary 4.2.
If $U_{\leq 3}$ is the set of generators for the invariants of degree at most three then $\operatorname{Frac}(\mathbb{C}[U_{\leq 3}])^{D_{n}}=\operatorname{Frac}(\mathbb{C}[x_{1},\ldots,x_{n}]^{D_{n}})$ and if $x\in Z(U_{\leq 3})^{c}$ then the orbit of $x$ is determined by the values of the degree three invariants.
Remark 4.3.
For this representation, the problem of recovering generic orbits of complex vectors from the unitary invariants of degree at most three is still open.
4.2. The complete multiplicity-free representation
From the perspective of MRA, a natural representation to consider is the sum of all irreducible representations but with each representation taken with multiplicity one. We call this representation the complete multiplicity-free representation. A finite group is abelian if and only every irreducible (complex) representation is one dimensional, if and only if the complete multiplicity-free representation equals the regular representation. Here we give examples to show that taking the sum of all irreducibles with multiplicity one is not sufficient to separate generic orbits.
These results show that the presence of irreducible representations with high multiplicity is an important feature of representations which have the property that invariants of low degree separate generic orbits. This is in contrast to degree bounds for full separating sets which, for finite groups, are independent of the multiplicities of the irreducibles [16].
4.2.1. The complete multiplicity-free representation of $D_{n}$
The irreducible representations of the dihedral group are as follows. If $n$ is even, then there are $(n/2-2)$ two-dimensional irreducible representations $V_{1},\ldots V_{(n/2-2)}$ where the rotation acts on $V_{\ell}$ with weights $e^{\pm 2\pi\iota\ell/n}$ and the reflection exchanges the two eigenspaces for the rotation. There are four characters, $L_{0},L_{-1},S_{0},S_{-1}$. The reflection acts trivially on $L_{0},L_{-1}$ and the rotation acts with weights 1 (i.e. trivially) and -1 respectively. On $S_{0},S_{-1}$ the reflection acts with weight $-1$ and the rotation acts with weights $1$ and $-1$ respectively.
If $n$ is odd then there are $(n-1)/2$ two-dimensional representations $V_{1},\ldots,V_{(n-1)/2}$ where the rotation acts on $V_{\ell}$ with eigenvalues $e^{\pm 2\pi\iota\ell/n}$ and the reflection exchanges the two eigenspaces. In addition if $n$ is odd $D_{n}$ has two characters, the trivial character $L_{0}$, and the character $S_{0}$ where the rotation acts trivially and the reflection acts with weight $-1$.
Note that the $n$-dimensional standard representation of $D_{n}$ where $D_{n}$ acts by cyclic shifts and reflection $x_{i}\mapsto x_{n-i}$ is the sum $L_{0}\oplus V_{1}+\ldots V_{n/2-2}\oplus L_{-1}$ if $n$ is even and $L_{0}\oplus V_{1}\ldots\oplus V_{(n-1)/2}$ is $n$ is odd. In particular it is multiplicity free.
Corollary 4.4.
Let $V$ be the complete multiplicity-free representation of the dihedral group $D_{n}$. Then the generic orbit $x\in V$ can only be recovered up to a list of size two for the invariants of degree at most three.
Remark 4.5.
Corollary 4.4 shows that adding new irreducibles to a multiplicity-free representation can increase the degrees of invariants needed to separate generic orbits.
Proof.
If $n$ is odd we can write an element of $V$ as $(x_{1},\ldots,x_{n},s_{0})$ where $(x_{1},\ldots,x_{n})$ are coordinates for the standard reprsentation and $s_{0}$ is coordinate for $S_{0}$. For a generic choice of $(x_{1},\ldots,x_{n})$ and non-zero $s_{0}$ the vectors $(x_{1},\ldots,x_{n},s_{0})$ and $(x_{1},\ldots,x_{n},-s_{0})$ lie in different $D_{n}$-orbits but their invariants of degree three or less are equal. In order to separate the orbits we need invariants of degree four. Precisely we need invariants of the form $s_{0}p(x_{1},\ldots,x_{n})$ where $p(x_{1},\ldots,x_{n})$ is invariant under rotations and the reflection acts by $-1$. However lowest degree of such a semi-invariant polynomial $p(x_{1},\ldots,x_{n})$ is three
If $n$ is even we can write an element of $V$ as $(x_{1},\ldots,x_{n},s_{0},s_{-1})$ and for a generic choice of $(x_{1},\ldots,x_{n})$ and $s_{0},s_{1}$ both non-zero, the vectors $(x_{1},\ldots,x_{n},s_{0},s_{-1})$ and $(x_{1},\ldots,x_{n},-s_{0},-s_{-1})$ have the same invariants of degree three or less but do not lie in the same $D_{n}$ orbit. ∎
4.2.2. The group $A_{4}$
The irreducible representations of $A_{4}$ are as follows: The three-dimensional irreducible representation $W$ which is the restriction of the standard irreducible representation $S_{4}$ and three characters $L_{0},L_{1}$ and $L_{2}$. By a standard result in representation theory, the one-dimensional representations are obtained as representations of the cyclic quotient $A_{4}/[A_{4},A_{4}]=\mathbb{Z}_{3}$. The character $L_{0}$ is the trivial representation, $L_{1}$ is the character where the 3-cycles act with weight $\omega=e^{2\pi\iota/3}$ and the products of disjoint transpositions act with weights $\omega^{2}$ and $L_{2}$ is the dual of $L_{1}$.
By Theorem 3.1 we know that for the regular representation $V=W^{3}+L_{0}+L_{1}+L_{2}$, the fraction field $\operatorname{Frac}(\mathbb{C}[V]^{A_{4}})$ is generated by invariant polynomials of degree at most three. However, the following proposition shows that if $V^{\prime}=W+L_{0}+L_{1}+L_{2}$ is the six-dimensional complete multiplicity-free representation of $A_{4}$ then invariants of degree at most three do not generate $\operatorname{Frac}(\mathbb{C}[V^{\prime}]^{A_{4}})$.
Proposition 4.6.
The subfield $K\subset\operatorname{Frac}(\mathbb{C}[V^{\prime}]^{A_{4}})$ generated by invariants of degree at most three has transcendence degree five over $\mathbb{C}$. In particular $\operatorname{Frac}(\mathbb{C}[V^{\prime}]^{A_{4}})$ is not generated by invariant polynomials of degree at most three.
Proof.
We start by making the action of $A_{4}$ on $\mathbb{C}[V^{\prime}]$ more explicit. Note that $W+L_{0}=S$ is the standard 4-dimensional permutation representation of $A_{4}$ and we use coordinates $(x_{1},\ldots,x_{4})$ on this summand where $\sigma\in A_{4}$ acts by permuting the coordinates. We denote by $x_{5},x_{6}$ the coordinates corresponding to the characters $L_{1}$ and $L_{2}$ respectively. By standard results in representation theory $\mathbb{C}[x_{1},\ldots,x_{4}]^{A_{4}}=\mathbb{C}[e_{1},e_{2},e_{3},e_{4},\Delta]$ where $e_{i}$ is the $i$-th elementary symmetric polynomial in $x_{1},x_{2},x_{3},x_{4}$ and $\Delta=(x_{4}-x_{3})(x_{4}-x_{2})(x_{4}-x_{1})(x_{3}-x_{2})(x_{3}-x_{1})(x_{2}-x_{1})$ is the Vandermonde polynomial. In particular the invariants of degree at most three generate a subfield of transcendence degree three.
The invariant polynomials in the last two coordinates are form $\mathbb{C}[x_{5},x_{6}]^{A_{4}}=\mathbb{C}[x_{5}^{3},x_{6}^{3},x_{5}x_{6}]$ and it is easy to check that the vector space of invariants of degree at most three $\mathbb{C}[x_{1},\ldots,x_{6}]^{A_{4}}_{\leq 3}$ is generated by the set $\{e_{1},e_{2},e_{3},x_{5}^{3},x_{6}^{3},x_{5}x_{6}\}$. These invariants clearly generate a subfield of $\operatorname{Frac}(\mathbb{C}[V^{\prime}]^{A_{4}})$ of transcendence degree 5 so they cannot separate generic orbits. ∎
Remark 4.7.
Proposition 4.6 implies that the invariants of degree at most three cannot list recover a generic vector in $V^{\prime}$. By Corollary 2.13 this implies that the moment tensors of degree at most three also cannot list recover a generic vector in $V^{\prime}$. In particular they cannot separate generic orbits.
4.3. The group $S_{n}$
Deep learning of objects such as sets or graphs which do not come with intrinsic ordering requires the construction of invariants which separate orbits under the symmetric group [3, 28, 30, 17]. Motivated by these questions, we consider the permutation action of $S_{n}$ on the space $\mathbb{C}^{n}\otimes(\mathbb{C}^{d})^{*}=\mathbb{C}^{n\times d}$ of $n\times d$ matrices, and pose the problem of determining bounds on the multiplicity $d$ which ensure that a set of very low-degree invariants can separate generic orbits.
Although $\mathbb{C}^{n\times d}$ is not strictly a subregular representation, it can be converted into one without any meaningful changes to the invariant theory. Recall that $\mathbb{C}^{n}$ splits as an $S_{n}$ representation into $\mathbb{C}^{n}=V_{0}+V$ where $V_{0}$ is the trivial representation generated by $(1,\ldots,1)$ and $V$ is its orthogonal complement which is the standard $(n-1)$-dimensional representation of $S_{n}$. Thus the space of matrices $(\mathbb{C}^{n})^{d}=V^{d}+V_{0}^{d}$ is not a subrepresentation of the regular representation when $d>1$, since the trivial representation only occurs with multiplicity one in the regular representation. However, we lose no important information by removing the trivial summand $V_{0}^{d}$ and working with just $V^{d}$ which is a subregular representation for $d\leq n-1$. Hence the space of matrices $\mathbb{C}^{n\times d}$ can appropriately be considered as a subregular representation for $d\leq n-1$.
When $d=1$ the ring of invariants $\mathbb{C}[\mathbb{C}^{n}]^{S_{n}}$ is a polynomial ring generated by the power sum symmetric polynomial (or alternatively the elementary symmetric polynomials) of degrees $1$ through $n$. Hence it is impossible to find a collection of invariants of degree less than $n$ that generate $\operatorname{Frac}(\mathbb{C}[\mathbb{C}^{n}]^{S_{n}})$.
When $d>1$ the ring of invariants is generated by multisymmetric polynomials of degrees $1$ through $n$ [27, 26]. As with the symmetric polynomials we have several different choices of generators for the ring of multisymmetric invariant polynomials. We will use the multisymmetric power sum polynomials. These polynomials are defined as follows: If $\{x_{i,j}\}$ for $1\leq i\leq n,1\leq j\leq d$ is a dual basis for $(\mathbb{C}^{n})^{d}$ then the sum $\sum_{\sigma\in S_{n},1\leq i_{1},...,i_{k}\leq d}\sigma(x_{1,i_{1}}...x_{1,i_{k}})=x_{1,i_{1}}...x_{1,i_{k}}+x_{2,i_{1}}...x_{2,i_{k}}+...+x_{n,i_{1}}...x_{n,i_{k}}$ is an $S_{n}$-invariant polynomial of degree $k$ (note that the $i_{1},...,i_{k}$ do not need to be distinct). We refer to invariants of this form as multisymmetric power sum polynomials.
When $d>1$ the power sum multisymmetric polynomials of degrees $1$ to $n$ are no longer algebraically independent and thus one can hope to find a set of invariants that separate generic orbits all of which have degree strictly less than $n$. Although this is an open problem for $n>3$, we give numerical evidence that for $d$ sufficiently large there is a set of invariants of degree at most three which form a transcendence basis for $\operatorname{Frac}(\mathbb{C}[V]^{S_{n}})$.
For $V=(\mathbb{C}^{n})^{d}$ there is a one-to-one correspondence between the power sum multisymmetric polynomials and the monomials of the form $x_{1,i_{1}}...x_{1,i_{k}}$ in the variables $\{x_{1,i}\}_{1\leq i\leq d}$. Therefore, the number of distinct power sum multisymmetric polynomials of degree $k$ is $\binom{d+k}{k}-1$. In particular there are $\binom{d+3}{3}+\binom{d+2}{2}+\binom{d+1}{1}-3=\frac{1}{6}(d^{3}+9d^{2}+26d)$ such invariants. Thus, for invariants of degree at most 3 to form a transcendence basis we must have that $\frac{1}{6}(d^{3}+9d^{2}+26d)\geq nd$. Asymptotically this forces $d$ to be at least $\sim\sqrt{n}$. In the table below we generated all of the multisymmetric power sum polynomials of degree at most three for different values of $n$ and $d\leq n-1$ and checked whether or not they form a transcendence basis using the Jacobian criterion as developed in [4, Proposition 3.25] and implemented in Mathematica.
Group Representation Invariants of degree at most three form a transcendence basis?
$S_{4}$ $\mathbb{C}^{4}$ No
$S_{4}$ $\mathbb{C}^{4\times 2}$ Yes
$S_{5}$ $\mathbb{C}^{5}$ No
$S_{5}$ $\mathbb{C}^{5\times 2}$ No
$S_{5}$ $\mathbb{C}^{5\times 3}$ Yes
$S_{6}$ $\mathbb{C}^{6}$ No
$S_{6}$ $\mathbb{C}^{6\times 2}$ No
$S_{6}$ $\mathbb{C}^{6\times 3}$ Yes
Table 1. Invariants and transcendence bases for symmetric groups
5. Proof of theorem 4.1 for real signals
In this section, we will work exclusively with real valued signals before extending the results to the complex case in the following section. To simplify many of the computations we identify $\mathbb{R}^{n}$ with its image in $\mathbb{C}^{n}$ under the Fourier transform. Thus we identify a vector $x\in\mathbb{R}^{n}$ with its Fourier transform $f=(f[0],f[1],\ldots,f[n-1])\in\mathbb{C}^{n}$ satisfying the condition that $f[i]=\overline{f[n-i]}$. Let $r$ and $s$ be the rotation and reflection which generate $D_{n}$. In the Fourier basis the action of $r$ can be diagonalized
(9) $r(f[\ell])=e^{\frac{2\pi\iota\ell}{n}}f[l]$
and the reflection action is given by
(10) $s(f[\ell])=f[n-\ell]=\overline{f{[\ell]}}.$
In the Fourier domain the representation $\mathbb{R}^{n}$ decomposes as a sum $L_{0}+V_{1}\ldots+V_{(n-1)/2}$ if $n$ is odd and $L_{0}+V_{1}\ldots+V_{n/2-1}+L_{-1}$ if $n$ is even. Here $L_{0}$ is the trivial representation, $L_{-1}$ the character where the rotation acts with weight $-1$ and reflection trivially. The representation $V_{\ell}$ is the two-dimensional representation of $D_{n}$ where $r$ acts with weights $e^{\pm 2\pi\iota\ell/n}$ and the reflection exchanges the eigenspaces.
The irreducible two-dimensional dihedral representation $V_{\ell}$ is the restriction of an irreducible two-dimensional $O(2)$ representation where the rotation $e^{\iota\theta}$ acts with weights $e^{\pm\iota\ell\theta}$ and the reflection exchanges the eigenspaces. Let $W_{k}=\oplus_{\ell=1}^{k}V_{\ell}$. We note that if we denote the trivial representation by $V_{0}$ then the representation $V_{0}+V_{1}+...+V_{k}$ is isomorphic to the space of $k$-band limited functions in $L^{2}(S^{1})$. Before addressing the dihedral case we analyze the $SO(2)$ and $O(2)$ orbit recovery problems for the representation $W_{k}$.
5.1. $SO(2)$ and $O(2)$-invariants
Theorem 5.1.
Let $W_{k}=\oplus_{\ell=1}^{k}V_{k}$ be the $2k$-dimensional real representation of $O(2)$.
  1. (1)
    The invariants of degree two and three for the restricted action of $SO(2)$ separate generic orbits in $W_{k}$.
  2. (2)
    If $k\neq 3$ the invariants of degree two and three separate generic orbits in $W_{k}$.
To establish notation let $e_{-k},\ldots,e_{k}$ be an orthonormal basis for $W_{k}$ in the Fourier domain and $e_{-\ell},e_{\ell}$ span $V_{k}$ where $e^{\iota\theta}\in SO(2)$ acts by $e^{\iota\theta}\cdot e_{m}=e^{\iota m\theta}e_{m}$. If we expand a vector $f\in W_{k}$ in terms of this basis as
(11) $f=\sum_{\ell=-k}^{k}f[\ell]e_{\ell}$
then the fact that $f$ is the Fourier transform of a real vector implies that $f[-\ell]=\overline{f[\ell]}$.
5.1.1. Proof for $SO(2)$-invariants
With this notation we have
(12) $T_{2}^{SO(2)}(f)=\sum_{i=1}^{k}{f}[i]{f}[-i]e_{i}e_{-i}=\sum_{i=1}^{k}|{f}[i]|^{2}e_{i}e_{-i}$
(13) $T_{3}^{SO(2)}(f)=\sum_{\{(i,j)|-k\leq i+j\leq k\}}{f}[i]{f}[j]{f}[-i-j]e_{i}e_{j}e_{-i-j}.$
Note that the degree-two invariant tensor determines the power spectrum (componentwise absolute value) of the signal $f$. We will show that if the Fourier coefficients $f[\ell]$ are non-vanishing for $1\leq\ell\leq k$ (and hence for $-k\leq\ell\leq-1$) then the degree-three invariant tensor determines a set of equations for the phases of the Fourier coefficients which determines them up to the action of $SO(2)$. Writing $f[j]=|{f}[j]|e^{i\theta_{j}}$ then, given the power spectrum, the third moment determines for all pairs $(i,j)$ with $i,j\geq 1$ and $i+j\leq k$
(14) $a_{ij}=e^{\theta_{i}+\theta_{j}-\theta_{i+j}}$
on the unit circle.
Proposition 5.2.
For a signal $f\in\mathbb{R}^{2k}$ with $f[i]\neq 0$ for $1\leq i\leq k$, the $a_{ij}$ defined in (11) determine the unknown phases up to translation by an element of $SO(2)$.
Proof.
From the $a_{ij}=\{e^{i(\theta_{i}+\theta_{j}-\theta_{i+j})}\}$ we can recursively express the phase $e^{i\theta_{\ell}}$ for $\ell>1$ in terms of $e^{i\theta_{1}}$. Precisely we have for $\ell\geq 2$
$e^{i\theta_{\ell}}=a_{1,\ell-1}^{-1}e^{i\theta_{1}}e^{i\theta_{\ell-1}}.$
5.1.2. $O(2)$ Invariants
Now for the orthogonal case. The degree-two invariant tensor
$T_{2}^{O(2)}(f)=\sum_{0\leq i\leq k}{f}[i]{f}[-i]e_{i}e_{-i}$
is the same as the degree-two $SO(2)$-invariant tensor and therefore gives us the power spectrum of $f$. The difference arises when we look at the third moment.
(15) $\begin{split}T_{3}^{O(2)}(f)&=\sum_{(i,j)\in S}({f}[i]{f}[j]{f}[-i-j]+{f}[-i]{f}[-j]{f}[i+j])(e_{i}e_{j}e_{-i-j}+e_{-i}e_{-j}e_{i+j})\\ &=\sum_{(i,j)\in S}({f}[i]{f}[j]{f}[-i-j]+\overline{{f}[i]{f}[j]{f}[-i-j]})(e_{i}e_{j}e_{-i-j}+e_{-i}e_{-j}e_{i+j})\end{split}$
where $S=\{(i,j)\in[1,k]^{2}|i+j\leq k\}$ Assuming all of the $f_{\ell}$ are non-vanishing the third moment then determines for each pair of indices $(i,j)$ with $i,j\geq 1$ and $i+j\leq k$
(16) $\alpha_{i,j}(f)=cos(\theta_{i}+\theta_{j}-\theta_{i+j})$
Since cosine is an even function, the quantity $\alpha_{i,j}$ determines the angle $(\theta_{i}+\theta_{j}-\theta_{i+j})$ up to sign or, equivalently the complex number $a_{i,j}(f)=e^{\iota(\theta_{i}+\theta_{j}-\theta_{i+j})}$ up to complex conjugation. Precisely, $a_{i,j}(f)=\alpha_{i,j}+i\sqrt{1-\alpha_{i,j}^{2}}$ or $a_{i,j}(f)=\alpha_{i,j}-i\sqrt{1-\alpha_{i,j}^{2}}$. The set of complex numbers $\{a_{i,j}(f)\}_{(i,j)}$ is invariant under rotation of $f$ by an element of $SO(2)$ and if $r\in O(2)\setminus SO(2)$ then $a_{i,j}(r\cdot f)=\overline{a_{i,j}(f)}$.
For a fixed vector $f$, consider the set of real vectors
$f^{\prime}=\sum_{\ell=1}^{k}f^{\prime}[-\ell]e_{-\ell}+f^{\prime}[\ell]e_{\ell}\in\mathbb{R}^{2k}$
that satisfy the equations
$T_{2}^{O(2)}(f^{\prime})=T_{2}^{O(2)}(f)$
and
$T_{3}^{O(2)}(f^{\prime})=T_{3}^{O(2)}(f).$
Since $f$ and $f^{\prime}$ have the same power spectrum we know that $f^{\prime}[\ell]=|f[\ell]|e^{\iota\varphi_{\ell}}$ for some angle $\varphi_{\ell}$. Equality of the degree-three invariants implies that for each pair of indices $(i,j)$ with $i+j\leq k$ $\alpha_{i,j}(f^{\prime})=\alpha_{i,j}(f)$ where $a_{i,j}(f^{\prime})=\cos(\varphi_{i}+\varphi_{j}-\varphi_{i+j})$. The next result shows that under a suitable genericity hypothesis on the Fourier coefficients of the original signal $f$, a single choice of the complex number $a_{p,q}(f^{\prime})=e^{\iota(\varphi_{p}+\varphi_{q}-\varphi_{p+q})}$ determines all other $a_{i,j}(f^{\prime})$ uniquely from the degree-three invariants. Because the signal $f^{\prime}$ is real we know that $a_{-i,-j}(f^{\prime})=\overline{a_{i,j}(f^{\prime})}$ so we need only determine the $a_{i,j}(f^{\prime})$ with $i,j\geq 1$. This in turn will imply that the set of $f^{\prime}\in\mathbb{R}^{n}$ which satisfy the equations $T_{2}^{O(2)}(f^{\prime})=T_{2}^{O(2)}(f)$ and $T_{3}^{O(2)}(f^{\prime})=T_{3}^{O(2)}(f)$ form a single $O(2)$ orbit.
Lemma 5.3.
The set of signals $f=\sum_{\ell=1}^{k}f[-\ell]e_{-\ell}+f[\ell]e_{\ell}$ with non-zero Fourier coefficients which satisfies the following condition
(*) There is no non-empty set of indices $S\subset\{(i,j)\subset[1,k]^{2}|i+j\leq k\}$ and all non-zero integers $n_{i,j}$ such that
(17) $\prod_{(i,j)\in S}\overline{a_{i,j}(f)}^{n_{i,j}}=\prod_{(i,j)\in S}a_{i,j}(f)^{n_{i,j}}$
is a non-empty $O(2)$-invariant Zariski open set in $\mathbb{R}^{2k}$
Proof.
We must show that the complement of the signals satisfying (*) is Zariski closed in the set of signals with non-zero Fourier coefficients. Given a non-empty set $S$ of indices if a signal $f$ satisfied (17) then multiplying both sides by $|f[i]f[j]f[i+j]|$ it also satisfies the non-zero real polynomial equation in its Fourier coefficients.
(18) $\prod_{(i,j)\in S}(f[i]f[j]f[-i-j])^{n_{i,j}}=\prod_{(i,j)\in S}(f[-i]f[-j]f[i+j])^{n_{i,j}}$
and hence lies in a proper real subvariety $X_{S}\subset\mathbb{R}^{2k}$. Thus, the set of vectors $f\in\mathbb{R}^{2k}$ with non-zero Fourier coefficients satisfying (*) is the Zariski open set $\mathbb{R}^{n}\setminus\left(\bigcup_{S}X_{S}\right)$. ∎
Proposition 5.4.
Suppose that for a pair of indices $(p,q)$ with $p+q\leq k$ we choose one of the two possible values for $a_{p,q}(f^{\prime})=e^{i(\varphi_{t}+\varphi_{p}-\varphi_{p+q})}$ then, if $k\geq 4$, all other $a_{i,j}(f^{\prime})$ are uniquely determined from the corresponding real numbers $\alpha_{i,j}$ provided that the signal $f$ satisfies condition (*) of Lemma 5.3
The proof of Proposition 5.4 requires the following combinatorial proposition.
Proposition 5.5.
If $k\geq 4$, then there exist integers $m_{i,j}$ all non-zero such that for any $(t_{1},\ldots,t_{k})\in\mathbb{R}^{k}$
$\sum_{\{(i,j)\geq 1|i+j\leq k\mathbb{\}}}m_{i,j}(t_{i}+t_{j}-t_{i+j})=0$
We now prove Proposition 5.4 assuming Proposition 5.5.
Proof of Proposition 5.4.
After possibly replacing $f$ with $r\circ f$ where $r\in O(2)$ we can assume $a_{p,q}(f^{\prime})=a_{p,q}(f)=\alpha$. Then we know that for every other pair of indices $(i,j)$ with $i+j\leq k$, $a_{i,j}(f^{\prime})=a_{i,j}(f)$ or $a_{i,j}(f^{\prime})=\overline{a_{i,j}}(f)$. Let $S$ be the set indices $(i,j)$ such that $a_{i,j}(f^{\prime})=\overline{\alpha_{i,j}(f)}$. By assumption $(p,q)\notin S$ so $S$ is a proper subset of the indices $(i,j)$. By Proposition 5.5 applied to the angles $\theta_{i}$ and $\varphi_{i}$ respectively we know that there are integers $m_{i,j}$ all non-zero such that
(19) $\prod_{(i,j)}a_{i,j}(f^{\prime})^{m_{i,j}}=\prod_{(i,j)}a_{i,j}(f)^{m_{i,j}}=1.$
Since $a_{i,j}(f^{\prime})=a_{i,j}(f)$ for $(i,j)\in S^{c}$ we can reduce (19) to
(20) $\prod_{(i,j)\in S}a_{i,j}(f^{\prime})=\prod_{(i,j)\in S}a_{i,j}(f).$
Since $a_{i,j}(f^{\prime})=\overline{a_{i,j}(f)}$ for $(i,j)\in S$ this means that $f$ does not satisfy hypothesis (*) unless $S=\emptyset$. Thus $\alpha_{i,j}(f^{\prime})=\alpha_{i,j}(f)$ for all pairs $(i,j)$. ∎
Proof of Proposition 5.5.
To prove Proposition 5.5 it suffices to prove if $x_{1},\ldots,x_{k}$ is the dual basis for $(\mathbb{R}^{k})^{*}$ then the forms $x_{ij}=\{x_{i}+x_{j}-x_{i+j}\}$ satisfy an equation of linear dependence
$\sum_{ij}q_{ij}x_{ij}$
with $q_{ij}\neq 0$ and rational for all pairs $(i,j)$. Since the $x_{ij}$ are integral linear combinations of the basis vectors $x_{i}$, we can view them as being elements of $(\mathbb{Q}^{k})^{*}$ and the result will follow from a few linear algebra lemmas.
Lemma 5.6.
Let $V$ be a vector space over an infinite field $K$ and let $v_{1},\ldots v_{m}$ be a collection of non-zero vectors in $V$ and let $W=\{(w_{1},...,w_{m})\}\subset K^{m}$ be the space of solutions to the equation $\sum_{i=1}^{m}w_{i}v_{i}=0$. Assume that for each $j$ there is a solution vector $(w_{1},...,w_{m})\in W$ with $w_{j}\neq 0$. Then there is a solution $(w_{1},...,w_{m})\in W$ with $w_{j}\neq 0$ for all $j$.
Proof.
By assumption $W$ is not contained in the hyperplane $V(x_{j})$ for each $i$. Hence, since $W$ is irreducible (because it is a linear subspace) it cannot be contained in the union of the $V(x_{j})$ which implies that there exists a solution vector $(w_{1},\ldots,w_{m})\in K^{m}$ with all $w_{j}$ non-zero. ∎
Lemma 5.7.
Let $V$ be a vector space over an infinite field $K$ and let $\{v_{1}...,v_{m}\}\subset V$ be a spanning set with $m>\dim V$. If for $i\leq m$ the set $S_{i}=\{v_{1}...,\hat{v_{i}},...,v_{m}\}$ still spans $V$ then there is an equation of linear dependence $\sum_{i=1}^{m}w_{i}v_{i}=0$ with $w_{i}\neq 0$ for all $i$.
Proof.
By Lemma 5.6 it suffices to find, for each $j$, a linear dependence $\sum_{i=1}^{k}w_{i}v_{i}=0$ with $w_{j}\neq 0$. By assumption $S_{j}$ spans $V$ so $v_{j}=\sum_{i\neq j}w_{i}v_{i}$ and hence the equation $\sum_{i=1}^{k}w_{i}v_{i}=0$ holds with $w_{j}=-1$. ∎
Definition 5.8.
We call a spanning set with the above property an excessive spanning set.
Proposition 5.5 now follows from the following lemma. ∎
Lemma 5.9.
Let $V^{-}_{k}$ be the hyperplane in $(\mathbb{Q}^{k})^{*}$ annihilated by the vector $(1,2,\ldots,k)\in\mathbb{Q}^{k}$. If $k\geq 2$ then the set of linear forms $\{x_{ij}=x_{i}+x_{j}-x_{i+j}\}\subset(\mathbb{Q}^{k})^{*}$ spans the hyperplane $V^{-}_{k}$ and if $k\geq 4$ it forms an excessive spanning set.
Proof.
First we note that $\langle x_{ij},(1,2,\ldots,k)\rangle=i+j-(i+j)=0$, so the span of the $x_{ij}$ lies in $V^{-}_{k}$. Let $L_{k}$ be the subspace of $(\mathbb{Q}^{k})^{*}$ spanned by the $\{x_{ij}\}$. We will prove by induction on $k$ that $L_{k}=V^{-}_{k}$. To start the induction note that if $k=2$ then $V^{-}_{2}$ is spanned by $x_{11}=2x_{1}-x_{2}$. To obtain the induction step observe that the embedding
$(\mathbb{Q}^{k})^{*}\subset(\mathbb{Q}^{k+1})^{*},(l_{1},\ldots,l_{k})\mapsto(l_{1},\ldots,l_{k},0)$
embeds $V^{-}_{k}$ into $V^{-}_{k+1}$. Assume by induction that the $V_{k}^{-}$ is spanned by $\{x_{ij}\}_{i+j\leq k}$; i.e. $V^{-}_{k}=L_{k}$. Then $L_{k+1}$ contains the additional linear form $x_{1}+x_{k}-x_{k+1}$ which is not in $L_{k}$. Hence $\dim L_{k+1}\geq\dim L_{k}+1=k$. On the other hand, we know that $L_{k+1}$ is contained in the hyperplane $V^{-}_{k+1}$. Therefore, $L_{k+1}=V^{-}_{k+1}$ since they are both hyperplanes.
A similar induction will allow us to show that if $k\geq 4$ then the $\{x_{ij}\}_{i+j\leq k}$ is an excessive spanning set of $V^{-}_{k}$. If $k=4$ then any three of the forms
$2x_{1}+x_{2},x_{1}+x_{2}+x_{3},x_{1}+x_{3}+x_{4},2x_{2}+x_{4}$
are linearly independent so these forms form an excessive spanning set for the hyperplane $V^{-}_{4}$. Assume by induction that $S_{k}=\{x_{ij}\}_{i+j\leq k}$ is an excessive spanning set for $V^{-}_{k}$. If we let $S_{k+1}=\{x_{ij}\}_{i+j\leq k+1}$ then $S_{k+1}=S_{k}\coprod T_{k+1}$ where $T_{k+1}=\{x_{ij}\}_{i+j=k+1}$. If we remove a vector $x_{ij}\in S_{k}\subset S_{k+1}$ then we know that $S_{k}\setminus\{x_{ij}\}$ spans $V^{-}_{k}$ by induction. Since none of the vectors in $T_{k}$ is contained in $V^{-}_{k}$ we conclude that $S_{k+1}\setminus\{x_{ij}\}$ spans all of $V^{-}_{k+1}$. On the other hand if $x_{i,k+1-i}$ is any vector in $T_{k+1}$ then $V^{-}_{k+1}$ is spanned by $S_{k}\cup\{x_{i,k+1-i}\}$ since $x_{i,k+1-i}=x_{i}+x_{k+1-i}-x_{k+1}$. Thus, as long as $T_{k+1}$ contains at least two vectors, which is the case if $k\geq 4$, then we may delete any vector from $S_{k+1}$ and still obtain a spanning set for $V^{-}_{k+1}$. ∎
Proof of Theorem 5.1 for $O(2)$.
We first consider the case where $k>3$. Given a signal $f$ satisfying the generic condition (*) of Lemma 5.3, suppose that $f^{\prime}$ has the same degree two and three invariant moments. Then $f$ and $f^{\prime}$ have the same power spectrum. If $k>3$ and $f$ satisfies condition (*) then Proposition 5.4 implies that after possibly replacing $f$ with $r\cdot f$ for some $r\in O(2)$ we have $a_{ij}(f^{\prime})=a_{ij}(f)$ for all pairs of indices $(i,j)$. By Proposition 5.2 this implies that $f^{\prime}$ and $f$ are in the same $SO(2)$ orbit after applying a reflection. Hence $f^{\prime}$ and $f$ are in the same $O(2)$ orbit.
If $k=1$ then $V_{1}$ is the defining representation of $O(2)$ and the second moment $f[-1]f[1]=|f[1]|^{2}$ uniquely determines the orbit. If $k=2$ and then the second moment of a vector $f$ determines the power spectrum of $f$ and $\cos(2\theta_{1}-\theta_{2})$. Rotating by an element of $SO(2)$ we can make $\theta_{1}$ arbitrary and for each choice of $\theta_{1}$ there are two possible values for $\theta_{2}$ which preserve the quantity $\cos(2\theta_{1}-\theta_{2})$. Hence, if $f$ has non-vanishing Fourier coefficients, the set of vectors with same degree two and three invariants consists of a single $O(2)$ orbit. ∎
Example 5.10.
When $k=3$ the invariants of degree at most three recover the $O(2)$ orbit of a generic vector up to a list of size two. If $f$ and $f^{\prime}$ have the same invariants of degree at most three than after possibly applying a reflection we can assume $a_{1,1}(f^{\prime})=a_{1,1}(f)$, and after applying a rotation that $e^{\iota\varphi_{1}}=e^{\iota\theta_{1}}$. Since $a_{1,1}(f^{\prime})=e^{2\iota\varphi_{1}+\varphi_{-2}}=a_{1,1}(f)=e^{2\iota\theta_{1}+\theta_{-2}}$. This is implies that $\varphi_{-2}=\theta_{-2}$, and hence that $\varphi_{2}=\theta_{2}$ since $f^{\prime}$ and $f$ are the Fourier transforms of real-valued functions. The only other piece of information we have is that
(21) $\cos(\varphi_{1}+\varphi_{2}-\varphi_{3})=\cos(\theta_{1}+\theta_{2}-\theta_{3}).$
However, for a fixed angle $\theta_{3}$ there are two angles $\varphi_{3}$ for which (21) holds.
Remark 5.11.
It is an interesting question as to the computational difficulty of recovering a signal from the orthogonal bispectrum. If we attempt to do frequency marching as in [8] we are forced to search over a set of $2^{n-1}$ conjugate possibilities before we determine the correct one. This is probably not optimal.
5.2. Recovering real vectors from cyclic invariants of low degree
The standard representation $\mathbb{C}^{n}$ of $\mathbb{Z}_{n}$ is the regular representation and hence by Corollary 3.1 we know that the invariants of degree at most three separate generic real and complex orbits. However, we will analyze this case directly as it will aid us in our proof that the dihedral invariants of degree at most three also separate generic dihedral orbits.
Case I. $n$ odd. If $n=2k+1$ is odd then the standard representation decomposes as $L_{0}+W_{k}$ where $W_{k}=\oplus_{\ell=1}^{k}V_{\ell}$ and $L_{0}$ is the trivial representation. If we expand $f\in\mathbb{R}^{n}$ as $f[0]f_{0}+\sum_{\ell=1}^{k}f[-\ell]e_{-\ell}+f[\ell]e_{\ell}$ with $f[-\ell]=\overline{f[\ell]}$ then the degree one invariant gives $f[0]$. We also know that the $SO(2)$-invariants (which are necessarily $\mathbb{Z}_{n}$ invariants) in degree three determine the $f[\ell]$ up to multiplication by $e^{\iota\ell\theta}$ for some angle $\theta$ which is independent of $\ell$. However if we choose $p,q,r>0$ so that $p+q+r=n$ then $f[p]f[q]f[r]$ is also a $\mathbb{Z}_{n}$-invariant which implies that $e^{\iota n\theta}=1$; i.e., the ambiguity is only up to to translation by an element of $\mathbb{Z}_{n}$.
Case II. $n$ is even. If $n=2k+2$ is even then the standard representation decomposes as $L_{0}+L_{-1}+W_{k}$, where $L_{0}$ is the trivial representation and $L_{-1}$ and is the character where the generator of $\mathbb{Z}_{n}$ acts by multiplication by $-1$. If we expand $f\in\mathbb{R}^{n}$ as $f=f[0]f_{0}+h[-1]f_{-1}+\sum_{\ell=1}^{k}f[-\ell]e_{-\ell}+f[\ell]e_{\ell}$ then once again the degree-one invariant is $f[0]$ and the same argument used in the previous case implies that the degree two and three invariants determine the Fourier coefficients $f[\ell]$ for $l=1,\ldots k$ up to the action of $\mathbb{Z}_{n}$. The Fourier coefficients $h[-1]$ is determined from the $f[\ell]$ and the degree-three $\mathbb{Z}_{n}$ invariant function $h[-1]f[1]f[n/2-1]$.
5.3. Dihedral invariants
We now follow the same approach for the dihedral invariants. The proofs for $n=2k+1$ and $n=2k+2$ are essentially identical as was the case for the cyclic invariants.
Case I. $n=2k+1$, $k\neq 3$ Suppose that $T_{d}^{D_{n}}(f)$ is known for $d\leq 3$. The Fourier coefficient $f[0]$ is once again known from the single degree-one invariant. By Theorem 5.1 we know that, after possibly replacing $f$ with its translate by the reflection $s\in D_{n}$, the subset of the $D_{n}$ invariants which are also $O(2)$ invariants determine the Fourier coefficients $f[\ell]$ up to multiplication by $e^{\iota\ell\theta}$ for some angle $\theta$. If $p+q+r=n$, then $D_{n}$-invariance of $f[p]f[q]f[r]+f[-p]f[-q]f[-r]$ implies that $e^{\iota n\theta}=1$, so $f$ is determined up to cyclic shift and rotation; i.e., by an element of $D_{n}$.
Case II. $n=2k+2$, $k\neq 3$ The same argument used in the odd case shows that the $D_{n}$ invariants of degree at most three determine the $f[0]$ and the Fourier coefficients $f[\ell]$ up to translation by an element of $D_{n}$. The coefficient $h[-1]$ is determined from the other Fourier coefficients by the degree-three invariant $h[-1](f[1]f[n/2-1]+f[n-1]f[n/2+1])$.
Because Theorem 5.1 requires $k\neq 3$ we need an additional proposition to cover the standard representations of the dihedral groups $D_{7}$ and $D_{8}$.
Proposition 5.12.
Let $\mathbb{R}^{7}$ and $\mathbb{R}^{8}$ be the standard representations of $D_{7}$ and $D_{8}$ respectively. The $D_{n}$-orbit of a generic vector $f\in\mathbb{R}^{n}$ is determined by the invariants of degree at most three.
Proof.
We will prove the result for $D_{7}$ but a similar argument will also work for $D_{8}$. Let $Z\subset\mathbb{R}^{7}$ be the $D_{7}$-invariant proper Zariski closed subset of vectors whose Fourier coefficients $f[\ell]$ satisfy the relation among $\mathbb{Z}_{7}$-invariants
(22) $\begin{split}(f[1]f[2]f[-3])^{\delta_{1}}(f[-2]^{2}f[-3])^{\delta_{2}}(f[1]f[3]^{2})^{\delta_{3}}=\\ (f[-1]f[-2]f[3])^{\delta_{1}}(f[2]^{2}f[3])^{\delta_{2}}(f[-1]f[-3]^{2})^{\delta_{3}}\end{split}$
for some $\delta_{1},\delta_{2},\delta_{3}\in\{0,1\}$ not all zero. We will prove that for any vector $f\in\mathbb{R}^{7}\setminus Z$ if $f^{\prime}\in\mathbb{R}^{7}$ satisfies $T_{d}^{D_{n}}(f^{\prime})=T_{d}^{D_{n}}(f)$ for $d=1,2,3$ then $f^{\prime}=g\cdot f$ for some $g\in D_{7}$.
Equality of the degree-one invariants immediately implies the equality of Fourier coefficients $f^{\prime}[0]=f[0]$ corresponding to the trivial summand in $\mathbb{R}^{7}$. Equality of the second moments implies that $f^{\prime}$ and $f$ have the same power spectra. Thus if $f[\ell]=r_{\ell}e^{\iota\theta_{\ell}}$ for some angles $\theta_{1},\theta_{2},\theta_{3}$ then $f^{\prime}[\ell]=r_{\ell}e^{\iota\varphi_{\ell}}$ for some angles $\varphi_{1},\varphi_{2},\varphi_{3}$.
Equality of the degree-three invariants
$\begin{array}[]{lll}f^{\prime}[1]^{2}f^{\prime}[-2]+f[-1]f^{\prime}[2]^{2}&=&f[1]^{2}f[-2]+f[-1]^{2}f[2]\\ f^{\prime}[1]f^{\prime}[2]f^{\prime}[-3]+f^{\prime}[-1]f^{\prime}[-2]f^{\prime}[3]&=&f[1]f[2]f[-3]+f[-1]f[-2]f[3]\\ f^{\prime}[1]f^{\prime}[3]^{2}+f^{\prime}[-1]f^{\prime}[-3]^{2}&=&f[1]f[3]^{2}+f[-1]f[-3]^{2}\\ f^{\prime}[2]^{2}f^{\prime}[3]+f^{\prime}[-2]^{2}f^{\prime}[-3]&=&f[2]^{2}f[3]+f[-2]^{2}f[-3]\end{array}$
implies that
$\alpha_{i,j}(f^{\prime})=\alpha_{i,j}(f)$
where $\alpha_{i,j}(f)=\cos(\theta_{i}+\theta_{j}-\theta_{i+j})$, $\alpha_{i,j}(f^{\prime})=\cos(\varphi_{i}+\varphi_{j}-\varphi_{i+j})$ and all indices are taken mod $\bmod\;7$. After possibly applying a reflection in $D_{7}$ we can assume that $a_{1,2}(f^{\prime})=a_{1,2}(f)=\alpha$ where $a_{i,j}(f)=e^{\iota(\theta_{i}+\theta_{j}-\theta_{i+j})}$, $a_{i,j}(f^{\prime})=e^{\iota(\varphi_{i}+\varphi_{j}-\varphi_{i+j})}$ and again all indices are taken $\bmod 7$.
The relation on the exponential expressions
$e^{\iota(\varphi_{1}+\varphi_{2}-\varphi_{3})}e^{-\iota(2\varphi_{2}+\varphi_{3})}e^{\iota(\varphi_{1}+2\varphi_{3})}=e^{\iota(2\varphi_{1}-\varphi_{2})}=\alpha$
implies that
(23) $a_{1,2}(f^{\prime})a_{-2,-2}^{2}(f^{\prime})a_{1,3}(f^{\prime})=\alpha=a_{1,2}(f)a_{-2,-2}^{2}(f)a_{1,3}(f).$
Let $S$ be the set of pairs $(i,j)$ such $a_{i,j}(f^{\prime})=\overline{a_{i,j}(f)}$. Then for $(i,j)\in S$ we have that $a_{i,j}(f^{\prime})=a_{-i,-j}(f)$. If $S$ is non-empty then after multiplying by the moduli of the Fourier coefficients we obtain an equation on the Fourier coefficients $f[\ell]$ of the form (22) with not all $\delta_{i}=0$. Thus if $f$ is generic then $S=\emptyset$; ie. $a_{i,j}(f^{\prime})=a_{i,j}(f)$. Hence by the reasoning of Section 5.2 we conclude that $f^{\prime}=r\cdot f$ for some rotation in $D_{7}$. Hence $f^{\prime}$ is obtained from $f$ after applying a possible reflection and rotation; i.e., by an element of $D_{7}$. ∎
6. Separation of complex dihedral orbits
Thus far we have proved that the $D_{n}$-orbit of a generic vector $x\in\mathbb{R}^{n}$ is determined by its first three invariant tensors. The argument made cruicial use of the fact that degree-two dihedral invariants determine the power spectrum of a real signal. This is no longer the case for signals $f\in\mathbb{C}^{n}$. To prove that generic complex orbits are separated by these invariants, we will use Proposition 2.15 and show that if $f\in\mathbb{R}^{n}$ is a generic signal then any complex signal $f^{\prime}$ with the same invariants of degree at most three must also be real.
Following our general proof strategy we will begin by proving this for cyclic invariants, even though this follows from Theorem 3.1.
6.1. Cyclic group
Lemma 6.1.
Let $f\in\mathbb{R}^{n}$ be a vector with all Fourier coefficients $f[\ell]$ non-vanishing. Suppose that $f^{\prime}\in\mathbb{C}^{n}$ satisfies $T^{\mathbb{Z}_{n}}_{d}(f^{\prime})=T^{\mathbb{Z}_{n}}_{d}(f)$ for $d\leq 3$ then $f^{\prime}\in\mathbb{R}^{n}$; i.e., $f^{\prime}[-i]=\overline{f^{\prime}[i]}$ for all $i$.
Proof.
First note that $f^{\prime}[0]=f[0]\in\mathbb{R}$ by assumption and if $n$ is even $f^{\prime}[n/2]^{2}=f[n/2]^{2}\in\mathbb{R}_{>0}$ which implies $f^{\prime}[n/2]\in\mathbb{R}$. In general we have $f^{\prime}[i]f^{\prime}[-i]=|f[i]|^{2}\in\mathbb{R}_{\geq 0}$ which implies that if $f^{\prime}[i]=r^{\prime}_{i}e^{\iota\varphi_{i}}$ then $f^{\prime}[-i]=r^{\prime}_{-i}e^{-\iota\varphi_{i}}$. It remains to show that $r^{\prime}_{i}=r^{\prime}_{-i}$ where the indices are taken $\bmod\;n$.
Since $f$ is a real vector we know that $f[-i]=\overline{f[i]}$ for all $i$. In particular for all pairs $(i,j)$ we have the equality of degree-three $\mathbb{Z}_{n}$ invariants
$f[i]f[j]f[-i-j]=\overline{f[-i]f[-j]f[i+j]}$
where the indices are taken $\bmod\;n$. Since $f^{\prime}$ and $f$ have the same degree-three $\mathbb{Z}_{n}$-invariants we can also conclude that
$f^{\prime}[i]f^{\prime}[j]f^{\prime}[-i-j]=\overline{f^{\prime}[-i]f^{\prime}[-j]f^{\prime}[i+j]}$
for all pairs $(i,j)$ with the indices taken $\bmod\;n$. This in turn implies that
(24) $r^{\prime}_{i}r^{\prime}_{j}r^{\prime}_{-i-j}=r^{\prime}_{-i}r^{\prime}_{-j}r^{\prime}_{i+j}.$
In particular $r^{\prime}_{2}=(\frac{r^{\prime}_{1}}{r^{\prime}_{-1}})^{2}r^{\prime}_{-2}$. Substituting this into the relation $r^{\prime}_{1}r^{\prime}_{2}r^{\prime}_{-3}=r^{\prime}_{-1}r^{\prime}_{-2}r^{\prime}_{-3}$ yields the relation $r^{\prime}_{3}=(\frac{r^{\prime}_{1}}{r^{\prime}_{-1}})^{3}r^{\prime}_{-3}$ and in general
(25) $r^{\prime}_{t}=\left(\frac{r^{\prime}_{1}}{r^{\prime}_{-1}}\right)^{t}r^{\prime}_{-t}.$
for $t\leq n/2$, so it suffices to prove that $r^{\prime}_{1}/r^{\prime}_{-1}=1$.
On the other hand if $i+j>n$, then substituting (25) into (24) yields
(26) $r^{\prime}_{i}r^{\prime}_{j}r^{\prime}_{n-i-j}=(\frac{r^{\prime}_{1}}{r^{\prime}_{-1}})^{i}r^{\prime}_{-i}(\frac{r^{\prime}_{1}}{r^{\prime}_{-1}})^{j}r^{\prime}_{-j}(\frac{r^{\prime}_{1}}{r^{\prime}_{-1}})^{n-i-j}r^{\prime}_{-n+i+j}$
(27) $=(\frac{r^{\prime}_{1}}{r^{\prime}_{-1}})^{n}r^{\prime}_{-i}r^{\prime}_{-j}r^{\prime}_{-n+i+j}=r^{\prime}_{-i}r^{\prime}_{-j}r^{\prime}_{-n+i+j}$
which yields $(\frac{r^{\prime}_{1}}{r^{\prime}_{-1}})^{n}=1$. Since we also know that $r_{1},r_{-1}$ are both positive we conclude that $r^{\prime}_{1}=r^{\prime}_{-1}$. ∎
We conclude the proof of Theorem 4.1 by proving a similar lemma for the dihedral group.
Lemma 6.2.
Let $f\in\mathbb{R}^{n}$ be a vector with all Fourier coefficients $f[\ell]$ non-vanishing. Suppose that $f^{\prime}\in\mathbb{C}^{n}$ satisfies $T^{D_{n}}_{d}(f^{\prime})=T^{D_{n}}_{d}(f)$ for $d\leq 3$ then $f^{\prime}\in\mathbb{R}^{n}$; i.e., $f^{\prime}[-i]=\overline{f^{\prime}[i]}$ for all $i$.
Proof.
The degree 2 invariants are the same as in the cyclic case and so we obtain ${f^{\prime}}[i]{f^{\prime}}[-i]\in\mathbb{R}$ implying that if $f^{\prime}[i]=r^{\prime}_{i}e^{i\varphi_{i}}$ then $f^{\prime}[-i]=r^{\prime}_{-i}e^{-i\varphi_{i}}$ $\forall i$. All that remains to show is that $r^{\prime}_{i}=r^{\prime}_{-i}$. As in the cyclic case, we have $f^{\prime}[0]\in\mathbb{R}$ and when $n$ is even $f^{\prime}[n/2]\in\mathbb{R}$ from the degree 1 and 2 invariants respectively.
The rest of the argument is similar to the cyclic case. For each pair $(i,j)$ we know that the dihedral invariant $f[i]f[j]f[-i-j]+f[-i]f[-j]f[-i-j]$ is real so
$f^{\prime}[i]f^{\prime}[j]f^{\prime}[-i-j]+f^{\prime}[-i]f^{\prime}[-j]f^{\prime}[i+j]=r_{i}r_{j}r_{-i-j}e^{\iota(\theta_{i}+\theta_{j}-\theta_{i+j})}+r_{-i}r_{-j}r_{i+j}e^{\iota(-\theta_{i}-\theta_{j}+\theta_{i+j})}$
is also real. Thus, $r^{\prime}_{i}r^{\prime}_{j}r^{\prime}_{-i-j}\sin(\varphi_{i}+\varphi_{j}-\varphi_{i+j})-r^{\prime}_{-i}r^{\prime}_{-j}r^{\prime}_{i+j}\sin(\varphi_{i}+\varphi_{j}-\varphi_{i+j})=0$ and hence $r^{\prime}_{i}r^{\prime}_{j}r^{\prime}_{-i-j}=r^{\prime}_{-i}r^{\prime}_{-j}r^{\prime}_{i+j}$ unless $\varphi_{i}+\varphi_{j}-\varphi_{i+j}$ is a multiple of $\pi$. However, if $\varphi_{i}+\varphi_{j}-\varphi_{i+j}$ is multiple of $\pi$ we can show that $r^{\prime}_{i}r^{\prime}_{j}r^{\prime}_{-i-j}=r^{\prime}_{-i}r^{\prime}_{-j}r^{\prime}_{i+j}$ with a different argument. Since $f$ is the Fourier transform of a real signal, then the equality of second moments $T_{2}(f^{\prime})=T_{2}(f)$ implies that $r^{\prime}_{\ell}r^{\prime}_{-\ell}=r_{\ell}^{2}$ for; i.e., $r^{\prime}_{-\ell}=r_{\ell}^{2}/r^{\prime}_{\ell}$. If $\sin(\varphi_{i}+\varphi_{j}-\varphi_{i+j})=0$ and we set $\beta=r^{\prime}_{i}r^{\prime}_{j}r^{\prime}_{-i-j}$ and $\alpha=r_{i}r_{j}r_{-i-j}$ then equality of degree-three invariants
$f^{\prime}[i]f^{\prime}[j]f^{\prime}[-i-j]+f^{\prime}[-i]f^{\prime}[-j]f^{\prime}[i+j]=f[i]f[j]f[-i-j]+f[-i]f[-j]f[-i-j]=2\alpha\cos(\theta_{i}+\theta_{j}-\theta_{i+j})$
implies that the non-negative real number $\beta$ satisfies the quadratic equation
(28) $\beta^{2}-2\beta\alpha\cos(\theta_{i}+\theta_{j}-\theta_{i+j})+\alpha^{2}=0$
which is impossible unless $\theta_{i}+\theta_{j}-\theta_{i+j}$ is also also multiple of $\pi$, since the discriminant of (28) is $4\alpha^{2}(\cos^{2}(\theta_{i}+\theta_{j}-\theta_{i+j})-1)$. In this case we must have $\beta=\alpha$ because they are both positive; i.e., $r^{\prime}_{i}r^{\prime}_{j}r^{\prime}_{-i-j}=r_{j}r_{j}r_{i-j}$. We also know that $\varphi_{-i}+\varphi_{-j}-\varphi_{-i-j}$ is also a multiple of $\pi$ so the same argument implies that $r^{\prime}_{-i}r^{\prime}_{-j}r^{\prime}_{i+j}=r_{-i}r_{-j}r_{i+j}=r_{i}r_{j}r_{-i-j}$, where the last equality holds because $f$ is real ∎
Example 6.3.
If some of the Fourier coefficients of a real signal $f$ vanish then there can be non-real signals $f^{\prime}$ with the same dihedral invariants of degree at most three. For example if $f\in\mathbb{R}^{5}$ has coordinates $(1,1,0,0,1)$ in the Fourier domain then the complex signal $f^{\prime}\in\mathbb{C}^{5}$ whose Fourier coordinates are $(1,1/2,0,0,2)$ has the same $D_{5}$ invariants of degrees at most three. In order to separate these orbits we need to use invariants of degree five.
In the course of our proof we have imposed various generic conditions on the signal $f$. However, by Corollary 4.2 we can give an explicit sufficient condition on the Fourier coefficients of a complex signal $f\in\mathbb{C}^{n}$ which guarantees that it can be recovered from its invariants of degree at most three.
Corollary 6.4.
If $f\in\mathbb{C}^{n}$ is a vector with the property that in the Fourier basis $f[i]\neq 0$ for all $i=0,\ldots,n-1$ and for each pair of indices $(i,j)$
$f[i]f[j]f[-i-j]+f[-i]f[-j]f[i+j]\neq 0$
then the $D_{n}$ orbit of $f$ is determined by the invariants of degree at most three.
Example 6.5.
The real vector $f$ whose coordinates in the Fourier basis are
$(f[0],f[1],f[2],f[3],f[4])=(1,\iota,-\iota,\iota,-\iota)$
has non-vanishing dihedral invariants of degree one and two. However, the polynomial invariants of degree three
$f[1]f[2]^{2}+f[4]f[3]^{2},f[1]^{2}f[3]+f[4]^{2}f[2]$
both vanish, and in this case the $D_{n}$ orbit is not uniquely determined by these invariants. Indeed the system of equations
(29) $\left\{\begin{array}[]{l}f^{\prime}[0]=1\\ f^{\prime}[1]f^{\prime}[4]=1\\ f^{\prime}[2]f^{\prime}[3]=1\\ f^{\prime}[1]^{2}f^{\prime}[3]+f^{\prime}[4]^{2}f^{\prime}[2]=0\\ f^{\prime}[1]f^{\prime}[2]^{2}+f^{\prime}[4]f^{\prime}[3]^{2}=0\end{array}\right.$
has twenty solutions corresponding to two real orbits of the dihedral group. Specifically the orbit of the real vector $f^{\prime}$ whose Fourier coordinates are $(1,-\iota,-\iota,\iota,\iota)$ has the same dihedral invariants but is not dihedrally equivalent to $f$. Note that $f^{\prime}$ does not have the same cyclic invariants since $f^{\prime}[1]f^{\prime}[2]^{2}=\iota$ but $f[1]f[2]^{2}=-\iota$